### v1

@@ 6.3 Generation and Understanding
Discussion and Implication: The analysis of the various subcategories within the ”Generation and Understanding” category reveals several insights into the challenges developers face and the implications for improving OpenAI’s models and their usage. (1) API Usage Issues. A significant portion of the challenges are related to the practical use of APIs. Issues such as repeated responses from the Davinci model, errors in embedding API usage, and problems with the Whisper API during audio processing are common. Developers often struggle with these API-related problems, seeking solutions and best practices for effective implementation. Therefore, detailed, updated documentation should be provided, especially for frequently encountered issues. For instance, recommending the use of the latest APIs, such as the Chat Completions API instead of older models like Davinci, can help mitigate repeated response issues. Providing best practices for embedding API usage, such as adding an offset vector to improve accuracy, can enhance developer experience. Similarly, sharing solutions for Whisper API errors, like addressing audio recording and encoding issues, can be beneficial. (2) Conceptual Challenges. Developers often face conceptual challenges, particularly with fine-tuning GPT models, parameter settings, iterative training, and output formatting, indicating a need for a deeper understanding of these processes. Therefore, creating comprehensive guides and tutorials on fine-tuning, common parameter settings, and iterative training techniques, can address many of these questions. (3) Application-Specific Issues. Challenges also arise from specific applications. Taking image processing as an example, developers encounter issues like translating prompts from other languages (such as French) to English, intellectual property concerns with generated images, and improving image recognition accuracy by adjusting contrast or rotation. Therefore, enhancing multilingual support for image processing prompts can help generate the desired images regardless of the input language. Providing clear guidelines on the intellectual property rights of generated images can address developers’ concerns and promote more confident use of the technology. Sharing techniques for improving image recognition accuracy, such as contrast adjustment and rotation correction, can help developers achieve better results in their applications.
6.4 Non-functional Properties
@@
The importance of non-functional properties cannot be ignored by OpenAI developers. These properties include aspects such as API call costs, rate limitations, and data security. In our taxonomy, 15.4% of the challenges belong to this category. The detailed subcategories are discussed as follows.
@@
Costs (D.1). Challenges in this category are related to API call costs. In the context of OpenAI development, costs are determined by the number of tokens utilized in each API call, particularly for generated content. Developers express concerns regarding the calculation of token usage and the cost [^46] https://community.openai.com/t/496993. They seek strategies to reduce costs by minimizing token consumption [^47] https://community.openai.com/t/525033. Responses to these challenges typically involve providing documentation-related information or offering strategies that consider factors such as the chosen model, input content length, and response content length. This category constitutes 3.6% of all challenges.
@@
Rate Limitation (D.2). Rate limitation is typically implemented to safeguard the stability and reliability of OpenAI services, preventing service interruptions or performance degradation due to excessive API usage. Developers need to understand the rate-limiting rules of the API, including the maximum number of calls per second, per minute, or per hour for different versions of the API, and how to manage call frequency in a sensible way to avoid triggering rate limits and causing service interruptions or exceptions. For example, developers encounter RateLimitError when calling gpt-3.5-turbo-0301484848https://community.openai.com/t/566696. Additionally, developers inquire whether rate limits are shared among different APIs [^49] https://community.openai.com/t/360331. Furthermore, developers ask about methods to increase the API rate limits [^50] https://community.openai.com/t/248374. These types of questions account for 3.2% of the total challenges.
@@
Regulation (D.3). When utilizing the OpenAI APIs, developers should strictly adhere to relevant regulations. For instance, OpenAI provides guidelines for the release of specific models or features and determines which users are granted permission to utilize them. Developers should review and abide by the terms of service and usage policies provided by OpenAI to ensure compliance with commercial and copyright regulations during API calls. For example, developers want to know whether OpenAI API can be used for commercial purposes [^51] https://community.openai.com/t/193738, or whether the images generated by DALL-E can be sold [^52] https://community.openai.com/t/33538. Additionally, developers encounter issues where prompts violate content policy rules [^53] https://community.openai.com/t/468967. This category constitutes 3.0% of all challenges.
@@
Promotion (D.4). Developers leverage forums as platforms to promote their developed plugins and GPTs, aiming to gather user feedback and potentially generate revenue. To this end, they offer comprehensive introductions to their creations, innovative functionalities, and usage guidelines. For instance, a developer promotes a plugin named “ShellMaster” [^54] https://community.openai.com/t/285122. Additionally, developers complain that the current GPT store lacks discoverability, meaning that many excellent GPTs are likely unknown to most people. They hope to add recommendations for users [^55] https://community.openai.com/t/585652. This category represents 2.1% of the challenges.
@@
Token Limitations (D.5). In addition to rate limiting, OpenAI imposes restrictions on the number of tokens for input and output during API calls to prevent abuse of API resources. Therefore, developers should understand the allocation, usage rules, and limitations of API tokens, as well as how to manage and optimize their usage. Specifically, developers find that the context length of the API is not fully utilized [^56] https://community.openai.com/t/291384. Additionally, developers seek methods to address token limitations  [^57] https://community.openai.com/t/4914. Furthermore, they inquire about solutions for handling large-scale datasets that exceed the token limitation and ineffective data splitting when using the gpt-4-1106-preview model [^58] https://community.openai.com/t/566336, as shown in Fig. 10. This category constitutes 2.0% of the challenges.
Figure 10: A sample post in token limitations (D.5)
Security (D.6). Ensuring security and preventing data leakage during API calls are the main concerns for OpenAI developers. They should prioritize protecting the privacy and integrity of data, minimizing the risk of data leaks due to security vulnerabilities or improper configurations. For example, developers express apprehension about potential data security risks when utilizing the API to process their uploaded files [^59] https://community.openai.com/t/354485. Additionally, they want to know whether using the API within internal networks can lead to data leakage [^60] https://community.openai.com/t/75725. This category constitutes 0.8% of all challenges.
@@
Privacy (D.7). Developers are advised to carefully review OpenAI’s privacy policy to gain insights into the data collection and processing practices during API calls and to understand the measures in place to protect users’ personal privacy. Furthermore, since API keys serve as critical qualifications for developer authentication and authorization, it is imperative for developers to prioritize the secure storage and management of these keys to prevent leaks or misuse. For example, developers create a desktop application based on ChatGPT and want to know if each user can use their API key within the program for billing purposes [^61] https://community.openai.com/t/161185. Additionally, developers have concerns regarding potential privacy issues associated with GPTs [^62] https://community.openai.com/t/496343. This subcategory constitutes 0.7% of all challenges.
@@
Discussion and Implication: Challenges such as API call costs, rate limitations, and token limitations are tightly linked to the development and usage of OpenAI’s services. Developers often express concerns about the costs associated with API calls, which are influenced by the choice of model and the number of tokens used in each request. Similarly, rate limitations are put in place to ensure service stability, but developers need to understand these limits and manage their API call frequencies accordingly. Token limitations pose additional challenges, especially when developers need to handle large datasets or require extensive context in their applications. Therefore, OpenAI should develop and provide tools that help developers accurately calculate and manage their token usage. This includes offering cost optimization strategies tailored to different models and usage scenarios. Moreover, providing detailed guidelines on model selection, including trade-offs between cost and performance, will help developers make informed decisions that align with their budget constraints.
General challenges such as safety and privacy are paramount when using AI services. Developers should ensure that their applications adhere to safety standards and protect user privacy. Therefore, OpenAI should continue to promote the use of its free moderation API, which helps reduce the frequency of unsafe content in completions. This tool should be integrated into the development process to automatically flag and filter potentially harmful outputs. Moreover, in high-stakes domains and code generation, it is crucial to have a human review of outputs before they are used in practice. This ensures that any limitations of the system are accounted for, and provides a safety net to verify the correctness and appropriateness of the generated content. Finally, limiting the amount of text a user can input into prompts and restricting the number of output tokens can help mitigate the risks of prompt injection and reduce the chances of misuse. These limitations should be implemented as part of best practices for safe AI deployment.
6.5 GPT Builder
@@
To meet specific user needs, OpenAI initially introduced the ChatGPT plugin, allowing developers to design plugins with various functionalities based on user requirements. Recently, OpenAI launched GPTs with similar functionality. However, due to their later release, there are fewer posts related to GPTs. Although OpenAI has deprecated plugins and recommends using GPTs as a replacement for users’ personalized needs, the OpenAI developer forum still lists plugins as an important category and combines plugins and GPTs into a single category. Moreover, plugin development and GPT development share many similarities in terms of development goals, technical foundations, and user usage. Therefore, we refer to the categories on the forum and collectively call both ChatGPT plugins and GPTs “GPT Builders” and analyze the related challenges.
@@
Development (E.1). In plugin development, developers encounter various challenges, including general and technical problems. Among general problems, developers seek guidance on selecting the most convenient IDE for plugin development [^63] https://community.openai.com/t/404743 and whether Markdown code can be integrated into plugin development [^64] https://community.openai.com/t/199399. Another category of challenges is related to technical issues. For instance, developers encounter a FatalServerError while developing a plugin [^65] https://community.openai.com/t/297501 or face parsing errors when attempting to connect GPTs to a cloud website [^66] https://community.openai.com/t/545528, as shown in Fig. 11. Additionally, developers ask how to use multiple GPTs simultaneously to handle the same task [^67] https://community.openai.com/t/552649. These challenges account for 11.2% of the total challenges.
Figure 11: A sample post in development (E.1)
@@
Testing (E.2). Following the development of plugins or GPTs, developers engage in software testing to ensure the correctness of the functionality implementation. During this process, developers encounter various issues. For instance, the openapi.yaml file might fail initial validation because it makes an HTTPS request instead of HTTP when testing the plugin on a local server [^68] https://community.openai.com/t/254680, or developers encounter errors while testing if the plugin is working correctly [^69] https://community.openai.com/t/166239. These challenges represent 0.9% of the total challenges.
@@
Discussion and Implication: Developers face numerous general challenges when developing plugins and GPTs, particularly in selecting the appropriate development environments and tools. This involves decisions about which Integrated Development Environments (IDEs) to use, which programming languages best suit their project needs, and how to set up their development environments efficiently. Therefore, OpenAI should develop comprehensive guides that cover the entire development lifecycle of plugins and GPTs, including recommendations for IDEs, programming languages, and other essential tools. This will help developers make informed choices that enhance their productivity and the quality of their work. Moreover, providing clear examples and templates can streamline the development process, particularly for those at varying skill levels. These resources can include step-by-step tutorials, sample projects, and best practices for developing and debugging, ensuring that developers have the support they need from project initiation to completion.
@@
After the development phase, rigorous testing is crucial to ensure the correct functioning of plugins and GPTs. Therefore, OpenAI should provide comprehensive automated testing tools that allow developers to test their plugins and GPTs efficiently. Automated testing can help identify bugs and issues early in the development cycle, reducing the time spent on manual testing and allowing for more consistent and repeatable tests. Moreover, these tools should support various testing methodologies, including unit tests, integration tests, and acceptance tests. By covering a broad range of scenarios, developers can ensure their applications are robust and reliable.
6.6 Prompt
@@
Prompts play a pivotal role in OpenAI API interactions, offering developers a mechanism to enhance the quality of the content produced. This category primarily focuses on challenges related to prompts, such as how to design and optimize them. Our findings indicate that 3.4% of the challenges fall into this category. Below, we analyze the specific subcategories it contains.
@@
Prompt Design (F.1). Developers often struggle with how to provide specific prompts to achieve the desired results and how to adjust their existing prompts. This includes techniques like targeting specific topics and managing the length of generated content to guide the model more accurately. For example, they seek advice on prompts to regenerate part of the output [^70] https://community.openai.com/t/214339(as shown in Fig. 12) and how to adjust prompts to improve the conversational style with the GPT model [^71] https://community.openai.com/t/301950. Challenges related to prompt design constitute 2.3% of the total questions.
Figure 12: A sample post in prompt design (F.1)
@@
Retrieval Augmented Generation (F.2). Retrieval Augmented Generation (RAG) Lewis et al (2020) is a technique that enhances the quality and relevance of generated text by integrating information retrieval with generation processes. This technique combines the strengths of retrieval and generation, enabling the model to better grasp the context and produce text that meets user expectations. Developers encounter issues where RAG fails to improve output quality [^72] https://community.openai.com/t/550286 or cause hallucinations after processing [^73] https://community.openai.com/t/408275. Questions also arise about preventing the RAG system from answering unrelated questions [^74] https://community.openai.com/t/434871. Challenges related to RAG account for 0.4% of the total challenges.
@@
Chain of Thought (F.3). The Chain of Thought (CoT) technique Wei et al (2022) is crucial for maintaining coherence and logical progression during text generation. It requires developers to craft input prompts that encourage the generation of text following a logical and semantic chain of thought. Achieving this involves careful selection of keywords, phrases, and sentence structures to steer the model toward producing coherent and structured text. Developers need a robust understanding of language, logical reasoning, and text generation capabilities, along with the skill to evaluate the model’s behavior and outputs. Ensuring semantic coherence and meeting expectations might prompt questions on effectively utilizing the CoT technique to construct prompts [^75] https://community.openai.com/t/128180 and seeking best practices for CoT in specific scenarios [^76] https://community.openai.com/t/17367. Challenges associated with this technique account for 0.2% of the total challenges.
@@
In-context Learning (F.4). In-context learning (ICL) (Dong et al, 2022) involves the model’s ability to learn and interpret based on the contextual information supplied within the input prompt. By meticulously designing these prompts, developers can embed specific contextual clues, enabling the model to deduce and produce content reflective of that context. For instance, there may be situations where, despite being provided with examples, the model repetitively generates identical outputs [^77] https://community.openai.com/t/17866. Developers seek advice on the recommended approach to provide context to follow-up questions to retrieve relevant context for completion [^78] https://community.openai.com/t/280666. Challenges related to in-context learning represent 0.2% of the total challenges.
@@
Zero-shot Prompting (F.5). Zero-shot prompting Kojima et al (2022) allows developers to leverage the model’s capabilities without needing sample data. Common issues include comparing the efficacy of zero-shot prompting versus fine-tuning for specific tasks [^79] https://community.openai.com/t/289714. Recently, some developers observed that zero-shot prompting works as expected on the OpenAI Playground but fails to generate valid responses when called via the Python API [^80] https://community.openai.com/t/216108. Challenges related to zero-shot prompting constitute 0.2% of the total questions.
@@
Tree of Thoughts (F.6). Tree of Thoughts (ToT) (Yao et al, 2024) is an advanced structured Chain of Thought (CoT) technique that guides the model toward generating text that is hierarchically organized and logically coherent. By creating detailed input prompts that integrate information from various branches and nodes within the ToT framework, developers enable the model to produce outputs with a clear hierarchy and logical structure. This technique assists the model in adhering to the ToT’s structure, facilitating the creation of more organized and logically structured outputs. Researchers have found that ToT significantly enhances language models’ problem-solving abilities in tasks requiring non-trivial planning or search, such as the Game of 24, creative writing, and mini crosswords [^81] https://community.openai.com/t/226512. Challenges related to ToT account for 0.1% of the total challenges.

### v3

@@ 6.3. Generation and Understanding
Discussion and Implication: The analysis of the various subcategories within the “Generation and Understanding” category reveals several insights into the challenges developers face and the implications for improving OpenAI’s LLMs and their usage. (1) API usage issues. A significant portion of the challenges are related to the practical use of APIs. Issues such as repeated responses from the LLM Davinci, errors in embedding API usage, and problems with the Whisper API during audio processing are common. Developers often struggle with these API-related problems, seeking solutions and best practices for effective implementation. Therefore, detailed, updated documentation should be provided, especially for frequently encountered issues. For instance, recommending the use of the latest APIs, such as the Chat Completions API instead of older models like Davinci, can help mitigate repeated response issues. Providing best practices for embedding API usage, such as adding an offset vector to improve accuracy, can enhance developer experience. Similarly, sharing solutions for Whisper API errors, like addressing audio recording and encoding issues, can be beneficial. Moreover, to implement API recommendations related to LLM development, a knowledge graph of relevant LLM APIs can be constructed based on API documentation and discussions from the OpenAI forum as suggested by Cao et al. (Cao et al., 2022). (2) Conceptual challenges. Developers often face conceptual challenges, particularly with fine-tuning GPT models, parameter settings, iterative training, and output formatting, indicating a need for a deeper understanding of these processes. Therefore, creating comprehensive guides and tutorials on fine-tuning, common parameter settings, and iterative training techniques, can address many of these questions. (3) Application-specific issues. Challenges also arise from specific applications. Taking image processing as an example, developers encounter issues like translating prompts from other languages (such as French) to English, intellectual property concerns with generated images, and improving image recognition accuracy by adjusting contrast or rotation. Therefore, enhancing multilingual support for image processing prompts can help generate the desired images regardless of the input language. Providing clear guidelines on the intellectual property rights of generated images can address developers’ concerns and promote more confident use of the technology. Sharing techniques for improving image recognition accuracy, such as contrast adjustment and rotation correction, can help developers achieve better results in their applications.
6.4. Non-functional Properties
@@
The importance of non-functional properties cannot be ignored by OpenAI developers and is closely intertwined with LLM development. These properties include aspects such as API call costs, rate limitation, and data security and privacy. In our taxonomy, 15.4% of the challenges belong to this category. The detailed subcategories are discussed as follows.
@@
Cost (D.1). Challenges in this subcategory are related to API call costs. In the context of OpenAI development, costs are determined by the number of tokens utilized in each API call, particularly for generated content. Developers express concerns regarding the calculation of token usage and the cost [^48] https://community.openai.com/t/496993. They seek strategies to reduce costs by minimizing token consumption [^49] https://community.openai.com/t/525033. Responses to these challenges typically involve providing documentation-related information or offering strategies that consider factors such as the chosen model, input content length, and response content length. This subcategory constitutes 3.6% of all challenges.
@@
Rate Limitation (D.2). Rate limitation is typically implemented to safeguard the stability and reliability of OpenAI services, preventing service interruptions or performance degradation due to excessive API usage. Developers need to understand the rate-limiting rules of the API, including the maximum number of calls per second, per minute, or per hour for different versions of the API, and how to manage call frequency in a sensible way to avoid triggering rate limits and causing service interruptions or exceptions. For example, developers encounter RateLimitError when calling gpt-3.5-turbo-0301505050https://community.openai.com/t/566696. Additionally, developers inquire whether rate limits are shared among different APIs [^51] https://community.openai.com/t/360331. Furthermore, as shown in Fig. 14, developers ask about methods to alleviate or increase the API rate limits [^52] https://community.openai.com/t/248374. These types of questions account for 3.2% of the total challenges.
Figure 14. A sample post in rate limitation (D.2).
@@
Regulation (D.3). When utilizing the LLM APIs, developers should strictly adhere to relevant regulations. For instance, the OpenAI organization provides guidelines for the release of specific models or features and determines which users are granted permission to utilize them [^53] https://openai.com/policies/usage-policies/. Developers should review and abide by the terms of service and usage policies provided by the OpenAI organization to ensure compliance with commercial and copyright regulations during API calls. For example, developers want to know whether LLM API can be used for commercial purposes [^54] https://community.openai.com/t/193738, or whether the images generated by the LLM DALL-E can be sold [^55] https://community.openai.com/t/33538. Additionally, developers encounter issues where prompts violate content policy rules [^56] https://community.openai.com/t/468967. This subcategory constitutes 3.0% of all challenges.
@@
Promotion (D.4). Developers often leverage forums as platforms to promote their developed plugins and GPTs, aiming to gather user feedback and potentially generate revenue. To this end, they offer comprehensive introductions to their creations, innovative functionalities, and usage guidelines. For instance, developers have expressed concerns about the lack of discoverability in the current GPT store, noting that many high-quality GPTs remain largely unknown to most users. To address this issue, they propose implementing a recommendation system to enhance user experience [^57] https://community.openai.com/t/585652. This category represents 2.1% of the challenges.
@@
Token Limitation (D.5). In addition to rate limiting, OpenAI imposes restrictions on the number of tokens for input and output during API calls to prevent abuse of API resources. Therefore, developers should understand the allocation, usage rules, and limitations of API tokens, as well as how to manage and optimize their usage. Specifically, developers find that the context length of the API is not fully utilized [^58] https://community.openai.com/t/291384. Additionally, developers seek methods to address token limitation  [^59] https://community.openai.com/t/4914. Furthermore, they inquire about solutions for handling large-scale datasets that exceed the token limitation and ineffective data splitting when using the LLM gpt-4-1106-preview [^60] https://community.openai.com/t/566336, as shown in Fig. 15. This category constitutes 2.0% of the challenges.
Figure 15. A sample post in token limitation (D.5).
@@
Security and Privacy (D.6). Ensuring data security and preventing leakage are primary concerns for OpenAI developers, especially when using API calls that involve sensitive information. Developers should focus on safeguarding data privacy, adhering to OpenAI’s privacy policies, and securely managing API keys to avoid unauthorized access and misuse. Specifically, first, ensuring security and preventing data leakage during API calls are the main concerns for OpenAI developers. For example, developers express apprehension about potential data security risks when utilizing the API to process their uploaded files [^61] https://community.openai.com/t/354485. Additionally, they want to know whether using the API within internal networks can lead to data leakage [^62] https://community.openai.com/t/75725. Second, developers are advised to carefully review OpenAI’s privacy policy to gain insights into the data collection and processing practices during API calls and to understand the measures in place to protect users’ personal privacy. Since API keys serve as critical qualifications for developer authentication and authorization, it is imperative for developers to prioritize the secure storage and management of these keys to prevent leaks or misuse. For example, developers create a desktop application based on ChatGPT and want to know if each user can use their API key within the program for billing purposes [^63] https://community.openai.com/t/161185. Finally, developers have concerns regarding potential privacy issues associated with GPTs [^64] https://community.openai.com/t/496343. This subcategory constitutes 1.5% of all challenges.
@@
Discussion and Implication: Challenges such as API call costs, rate limitation, and token limitation are tightly linked to the development and usage of the LLM’s services. Developers often express concerns about the costs associated with API calls, which are influenced by the choice of model and the number of tokens used in each request. Similarly, rate limitations are put in place to ensure service stability, but developers need to understand these limits and manage their API call frequencies accordingly. Token limitations pose additional challenges, especially when developers need to handle large-scale datasets or require extensive context in their applications. Therefore, the LLM vendors should develop and provide tools that help developers accurately calculate and manage their token usage. These tools should further offer cost optimization strategies (Chen et al., 2023c) tailored to different models and usage scenarios. Moreover, providing detailed guidelines on model selection, including trade-offs between cost and performance, will help developers make informed decisions that align with their budget constraints.
General challenges such as safety and privacy are paramount when using AI services. For LLM vendors, it is imperative to ensure compliance with global data protection regulations, such as GDPR (General Data Protection Regulation). This involves providing more transparent data processing and model training workflows to meet regulatory requirements. Additionally, LLMs may become targets of malicious attacks, such as prompt injection (Liu et al., 2023), where the model output is manipulated to generate harmful or misleading content. Therefore, researching and implementing effective defense mechanisms is critical to safeguard their applications. For LLM developers, especially in high-stakes domains (such as healthcare, finance, and software development), manual review of model outputs before their practical use is essential to mitigate risks caused by hallucinations (Ahmad et al., 2023; Kang and Liu, 2023; Liu et al., 2024b; Zhang et al., 2024). Furthermore, after constructing a prompt, developers should utilize OpenAI’s Moderations API [^65] https://platform.openai.com/docs/api-reference/moderations to assess whether the prompt contains harmful content, ensuring the safety and integrity of the generated outputs.
6.5. GPT Builder
@@
To meet specific user needs, the OpenAI organization initially introduced the ChatGPT plugin, allowing developers to design plugins with various functionalities based on user requirements. Recently, OpenAI launched GPTs with similar functionality. However, due to their later release, there are fewer posts related to GPTs. Although OpenAI has deprecated plugins and recommends using GPTs as a replacement for users’ personalized needs, the OpenAI developer forum still lists plugins as an important subforum and combines plugins and GPTs into a single subforum. Moreover, plugin development and GPT development share many similarities in terms of development goals, technical foundations, and user usage. Therefore, we call both ChatGPT plugins and GPTs “GPT Builder” and analyze the related challenges.
@@
Development (E.1). In plugin development, developers encounter various challenges, including general and technical problems. Among general problems, developers seek guidance on selecting the most convenient IDE for plugin development [^66] https://community.openai.com/t/404743 and whether Markdown can be integrated into plugin development [^67] https://community.openai.com/t/199399. Another category of challenges is related to technical issues. For instance, developers encounter a FatalServerError while developing a plugin [^68] https://community.openai.com/t/297501 or face parsing errors when attempting to connect GPTs to a cloud server [^69] https://community.openai.com/t/545528, as shown in Fig. 16. Additionally, developers ask how to use multiple GPTs simultaneously to handle the same task [^70] https://community.openai.com/t/552649. These challenges account for 11.2% of the total challenges.
Figure 16. A sample post in development (E.1).
@@
Testing (E.2). After developing plugins or GPTs, developers conduct software testing to verify the correctness of the implemented functionalities. Throughout this process, they often encounter a range of challenges. For instance, the openapi.yaml file might fail initial validation because it makes an HTTPS request instead of HTTP when testing the plugin on a local server [^71] https://community.openai.com/t/254680, or developers encounter errors while testing if the plugin is working correctly [^72] https://community.openai.com/t/166239. In addition, some functional errors may be encountered during testing. For example, when developers test a GPT action, they find that after clicking to test, only the name of the GPT is displayed, without any other content output [^73] https://community.openai.com/t/597697/6. These challenges represent 0.9% of the total challenges.
@@
Discussion and Implication: Developers face numerous general challenges when developing plugins and GPTs, particularly in selecting the appropriate development environments and tools. This involves decisions about which Integrated Development Environments (IDEs) to use, which programming languages best suit their project needs, and how to set up their development environments efficiently. Therefore, the OpenAI organization should provide comprehensive guides that cover the entire development lifecycle of plugins and GPTs, including recommendations for IDEs, programming languages, and other essential tools. This will help developers make informed choices that enhance their productivity and the quality of their developed products. Moreover, providing more presentative examples and templates can streamline the development process, particularly for those at varying skill levels. These resources can include step-by-step tutorials, sample projects, and best practices for developing and debugging, ensuring that developers have the support they need from project initiation to completion.
@@
Following the development phase, rigorous testing is essential to ensure the proper functionality of plugins and GPTs. Therefore, the OpenAI organization should provide comprehensive automated testing tools that allow developers to test their plugins and GPTs efficiently. Automated testing can help identify bugs early in the development cycle, reducing the time spent on manual testing. Moreover, these testing tools should support multiple phases of the testing process, including unit testing, integration testing, and acceptance testing. By encompassing a comprehensive range of phases, developers can ensure their applications are both robust and reliable.
6.6. Prompt
@@
Prompts play a pivotal role in LLM interactions, offering developers a mechanism to enhance the quality of the content produced. This category primarily focuses on challenges related to prompts, such as how to design and optimize them. Our findings indicate that 3.4% of the challenges fall into this category. The detailed subcategories are discussed as follows.
@@
Prompt Design (F.1). Developers often struggle with how to provide specific prompts to achieve the desired results and how to modify their existing prompts. This includes techniques like targeting specific topics and managing the length of generated content to guide the model more accurately. For example, they seek advice on prompts to regenerate part of the output [^74] https://community.openai.com/t/214339, which is shown in Fig. 17, and how to modify prompts to improve the conversational style with the GPT model [^75] https://community.openai.com/t/301950. Challenges related to prompt design constitute 2.3% of the total challenges.
Figure 17. A sample post in prompt design (F.1).
@@
Retrieval Augmented Generation (F.2). Retrieval Augmented Generation (RAG) (Lewis et al., 2020) is a technique that enhances the quality and relevance of generated text by integrating information retrieval with generation processes. This technique combines the strengths of retrieval and generation, enabling the model to better grasp the context and produce text that meets user expectations. Developers encounter issues where RAG fails to improve output quality [^76] https://community.openai.com/t/550286 or cause hallucinations after processing [^77] https://community.openai.com/t/408275. Questions also arise about preventing the RAG system from answering unrelated questions [^78] https://community.openai.com/t/434871. Challenges related to RAG account for 0.4% of the total challenges.
@@
Chain of Thought (F.3). The Chain of Thought (CoT) technique (Wei et al., 2022) is crucial for maintaining coherence and logical progression during text generation. It requires developers to craft input prompts that encourage the generation of text following a logical and semantic chain of thought. Achieving this involves careful selection of keywords, phrases, and sentence structures to steer the model toward producing coherent and structured text. Developers need a robust understanding of language, logical reasoning, and text generation capabilities, along with the skill to evaluate the model’s behavior and outputs. Ensuring semantic coherence and meeting expectations might prompt questions on effectively utilizing the CoT technique to construct prompts [^79] https://community.openai.com/t/128180 and seeking best practices for CoT in specific scenarios [^80] https://community.openai.com/t/17367. Challenges associated with the COT technique account for 0.2% of the total challenges.
@@
In-context Learning (F.4). In-context learning (ICL) (Dong et al., 2022) involves the LLM’s ability to learn and interpret based on the contextual information supplied within the input prompt. By meticulously designing these prompts, developers can embed specific contextual clues, enabling the model to deduce and produce content reflective of that context. For instance, there may be situations where, despite being provided with examples, the model repetitively generates identical outputs [^81] https://community.openai.com/t/17866. Developers seek advice on the recommended approach to provide context to follow-up questions to retrieve relevant context for completion [^82] https://community.openai.com/t/280666, as shown in Fig. 18. Challenges related to in-context learning represent 0.2% of the total challenges.
Figure 18. A sample post in in-context learning (F.4).
@@
Zero-shot Prompting (F.5). Zero-shot prompting (Kojima et al., 2022) allows developers to leverage the model’s capabilities without needing sample data. Common issues include comparing the efficacy of zero-shot prompting versus fine-tuning for specific tasks [^83] https://community.openai.com/t/289714. Recently, some developers observed that zero-shot prompting works as expected on the OpenAI Playground but fails to generate valid responses when called via the Python API [^84] https://community.openai.com/t/216108. Challenges related to zero-shot prompting constitute 0.2% of the total challenges.
@@
Tree of Thoughts (F.6). Tree of Thoughts (ToT) (Yao et al., 2024) is an advanced structured Chain of Thought (CoT) technique that guides the model toward generating text that is hierarchically organized and logically coherent. By creating detailed input prompts that integrate information from various branches and nodes within the ToT framework, developers enable the model to produce outputs with a clear hierarchy and logical structure. This technique assists the model in adhering to the ToT’s structure, facilitating the creation of more organized and logically structured outputs. Researchers have found that ToT significantly enhances language models’ problem-solving abilities in tasks requiring non-trivial planning or search, such as the Game of 24, creative writing, and mini crosswords [^85] https://community.openai.com/t/226512. Challenges related to the tree of thoughts account for 0.1% of the total challenges.